# Deep Learning assisted Port-Cycling based Channel Sounding for Precoder Estimation in Massive MIMO Arrays

### v1

@@ Abstract
Future wireless systems are expected to employ a substantially larger number of transmit ports for channel state information (CSI) estimation compared to current specifications. Although scaling ports improves spectral efficiency, it also increases the resource overhead to transmit reference signals across the time-frequency grid, ultimately reducing achievable data throughput. In this work, we propose an deep learning (DL)-based CSI reconstruction framework that serves as an enabler for reliable CSI acquisition in future 6G systems. The proposed solution involves designing a port-cycling mechanism that sequentially sounds different portions of CSI ports across time, thereby lowering the overhead while preserving channel observability. The proposed CSI Adaptive Network (CsiAdaNet) model exploits the resulting sparse measurements and captures both spatial and temporal correlations to accurately reconstruct the full-port CSI. The simulation results show that our method achieves overhead reduction while maintaining high CSI reconstruction accuracy.
@@ I Introduction
The massive multiple-input multiple-output (MIMO) system is a key technology for wireless communication systems, offering benefits including spatial multiplexing, capacity gain, etc. To achieve these benefits, accurate CSI is vital, which incurs substantial reference signal overhead with increasing channel dimensions. In current systems, downlink channel acquisition is performed using CSI-RS [1]. It is useful in assessing channel conditions and reports back the channel quality, the number of independent streams, and the beamforming vector to be utilized [2]; however, its overhead grows linearly with the array size thereby reducing the resources available for data transmission.
To address the growing overhead, industry and academic efforts have explored Artificial Intelligence/ Machine Learning (AI/ML) techniques for CSI feedback. In particular, 3GPP has identified AI/ML based CSI compression and prediction as initial use-cases for AI/ML based air interface [3]. From an academic perspective, CsiNet was the pioneer in demonstrating effective non-linear compression of CSI with reduced payload sizes [4]. In [5], authors have investigated CSI extrapolation across time, frequency and spatial domains to actively work on feedback reduction. While these approaches reduce feedback overhead or solve temporal issues, they do not address the instantaneous overhead associated with sounding a large number of antenna ports, which remains a main bottleneck in massive MIMO systems.
@@
Building on this framework, we propose CsiAdaNet, a deep learning based CSI reconstruction model that exploits the temporal and spatial correlations across the temporally aggregated CSI measurements. By coherently aggregating these sparse observations across time, CsiAdaNet predicts the quantized Type-II precoder defined in 5G NR [1],[7],enabling precoder estimation without requiring full-port instantaneous CSI.
@@ II System Model
where $\mathbf{y}_{k}\in\mathbb{C}^{N_{r}\times 1}$ denotes the received signal, $\mathbf{h}_{k}\in\mathbb{C}^{N_{r}\times N_{t}}$ denotes the channel matrix, $\mathbf{v}_{k}\in\mathbb{C}^{N_{L}\times N_{t}}$ denotes the precoding vector, $\mathbf{x}_{k}\in\mathbb{C}^{N_{L}\times 1}$ denotes the data vector and $n_{k}$ denotes the additive noise of the $k^{th}$ subcarrier. The BS is equipped with $N_{t}=2\times N_{x}\times N_{y}$ antennas, where $N_{x},N_{y}$ denotes the number of dual polarized antenna elements at the horizontal and vertical dimensions. The channel matrix $\mathcal{H}\in\mathbb{C}^{N_{r}\times N_{t}\times K}$ is obtained by stacking the channel matrices of each CSI subcarrier respectively.
@@ II-A Type II codebook in 5G NR
  • Wideband Amplitude: It corresponds to reporting the quantized wideband amplitudes for the selected 2L12L-1 orthogonal beams. The strongest beam within the beam set is normalized to unit amplitude. The quantized amplitudes are drawn from the set, $P={\{0,\sqrt{\frac{1}{64}},\sqrt{\frac{1}{32}},\sqrt{\frac{1}{16}},\sqrt{\frac{1}{8}},\sqrt{\frac{1}{4}},\sqrt{\frac{1}{2}},1\}}$
@@
1:Input: Over-sampled beam grid $\{\mathbf{V}_{b}\}_{b=1}^{B}$, covariance $\mathbf{R}\in\mathbb{C}^{N_{t}\times N_{t}}$, $L$ : Number of beams
2:Output: Beam Set ($B^{\star}$), Beam Indices ($\mathcal{I}^{\star}$)
3:
$\mathbf{R}=\begin{bmatrix}\mathbf{R}_{11}&\mathbf{R}_{12}\\ \mathbf{R}_{21}&\mathbf{R}_{22}\end{bmatrix}$
4:for $b=1$ to $B$ do
5:  $\mathbf{r}\leftarrow\mathbf{V}_{b}^{H}(\mathbf{R}_{11}+\mathbf{R}_{22})\mathbf{V}_{b}$
6:  $\mathbf{s}\leftarrow|\operatorname{diag}(\mathbf{r})|$
7:  $\mathcal{I}_{b}\leftarrow\operatorname{TopL}(\mathbf{s})$
8:  $\eta_{b}\leftarrow\sum_{i\in\mathcal{I}_{b}}s_{i}$
9:end for
@@
11:return $B^{\star},\mathcal{I}^{\star}$
@@
1:Input: Beamset $b^{\star}$, beam indices $\mathcal{I}^{\star}$, covariance matrix $\mathbf{R}\in\mathbb{C}^{N_{t}\times N_{t}}$
2:Output: Wideband Amplitude ($\mathcal{P}_{\nu}^{\star}$), Wideband Phase ($\mathcal{\varphi}_{\nu}^{\star}$)
@@
5:for $l=1$ to $L$ do
6:  $p^{\nu}_{0,l}\leftarrow\mathbf{e}_{0,\nu}^{H}\mathbf{v}_{l}$, $p^{\nu}_{1,l}\leftarrow\mathbf{e}_{1,\nu}^{H}\mathbf{v}_{l}$
7:end for
8:for $\text{pol}\in\{0,1\}$ do
9:  $a^{\nu}_{pol}\leftarrow|p^{\nu}_{pol}|$
10:  $k^{\star}_{\nu,pol}\leftarrow\arg\max_{k}a^{\nu}_{k,pol}$
11:  $\tilde{{a}^{\nu}_{pol}}\leftarrow{a}^{\nu}_{pol}/a_{k^{\star}_{\nu,pol}}$
12:  $\mathbf{A}_{\nu,pol}\leftarrow|\tilde{\mathbf{a}^{\nu}_{pol}}|,\quad\boldsymbol{\theta}_{\nu,pol}\leftarrow\angle\tilde{\mathbf{a}^{\nu}_{pol}}$
13:  for $l=1$ to $L$ do
14:   $\mathcal{A}^{\nu}_{pol,l}\leftarrow\arg\min_{i}\left|P_{i}-{A}^{\nu}_{pol,l}\right|$
15:   ${\Theta}^{\nu}_{pol,l}\leftarrow\arg\min_{j}\left|\Phi_{j}-{\theta}^{\nu}_{pol,l}\right|$
16:  end for
17:end for
[…]
19:return $\mathcal{P}_{\nu}^{\star}$, $\mathcal{\varphi}_{\nu}^{\star}$
Algorithm 3 Sub-panel Partitioning
1:Input: UPA size $(N_{x},N_{y})$, Number of partitions $\rho=(\rho_{x}\times\rho_{y})$
2:Output: Set of sub-arrays $\mathcal{S}$
[…]
5:for $i=0$ to $\rho_{x}-1$ do
6:  for $j=0$ to $\rho_{y}-1$ do
7:   $\mathcal{S}_{i,j}\leftarrow$ antenna elements in rows $iN_{x}^{\prime}:(i{+}1)N_{x}^{\prime})]$ and cols $jN_{y}^{\prime}:(j{+}1)N_{y}^{\prime})]$
8:   $\mathcal{S}\leftarrow\mathcal{S}\cup\{\mathcal{S}_{i,j}\}$
9:  end for
10:end for
11:return $\mathcal{S}$
@@ II-B Sub-array partitioning
The main motivation in this work is to utilize a lesser number of ports for active transmission of reference signals. To reduce the number of active ports for transmission, the full-planar array is partitioned into smaller sub-arrays. In general, the active subset of ports can be selected using sub-panel, interleaved or arbitrary configurations. In this work, we opt to model the sub-array based on sub-panel partitioning, as it preserves the antenna spacing, thereby avoiding spatial aliasing in the beam computation.
The uniform planar array of size ($N_{x}\times N_{y}$) is divided into $\rho=\rho_{x}\times\rho_{y}$ contiguous sub-panels, each consisting of $N^{\prime}_{x}=N_{x}/\rho_{x}$ and $N^{\prime}_{y}=N_{y}/\rho_{y}$ antenna elements. This results in $\rho$ non-overlapping sub-arrays, each containing $N^{\prime}_{t}=N^{\prime}_{x}\times N^{\prime}_{y}\times N_{pol}$ ports as mentioned in Algorithm 3. and an example configuration is illustrated in Fig. 1.
@@ II-C Port Cycling Scheme
It is a method of cycling through the available subset of ports at different time instants for downlink reference signal transmission. The motivation is that, when beam related quantities such as beamset ($\mathcal{B}$), dominant beam index ($\mathcal{I}_{0}$) do not vary across each of the sub-array along time, it underlines the fact that wireless channel statistics such as dominant path, angular spread remain relatively stable within the transmission duration. A score is defined to measure the degree of variation of each of the beam related metric ($\mathcal{M}$).
@@
Figure 2: Plot of variation scores
As shown in Fig. 2, both the beamset index and dominant beam index exhibit low variation across port-cycling instances, particularly at moderate-to-high SNR region. The beamset remains highly stable even under noisy conditions, while the beam index converges quickly with increasing SNR. These observations validate the feasibility of temporally aggregating port-cycled CSI measurements to predict precoder.
@@
III CsiAdaNet
This section provides details about a Deep Learning assisted solution for the port cycling based system. The proposed CSIAdaNet is an DL model that resides at the user equipment (UE) side and operates on $\rho$ periodic CSI measurements, each corresponding to a distinct sub-panel based sub-array with a periodicity of $P$ ms. After cycling through all the $\rho$ sub-panels and estimating the channel corresponding to them, the model predicts the quantized Type-II codebook parameters.
$B^{\star},\;\mathcal{I}^{\star},\;\mathcal{P}^{\star},\mathcal{\varphi}^{\star}=\mathcal{F}(\{\mathcal{H}_{t+iP}\}_{i=0}^{\rho-1};\theta)$ (3)
where $\mathcal{F}$ denotes the DL model, $\theta$ refers to the model parameters and $\mathcal{H}_{t+iP}$ denoting the CSI corresponding to sub-array sounded at time $t+iP$.
The model performs the predictions sequentially with contextual dependency. The beamset is initially predicted, followed by the prediction of the top $L$ beam indices conditioned on the beamset. The predicted beamset and beam indices are then jointly used to estimate the quantized wideband amplitude levels and phase levels, enabling a hierarchical passage of information that remains consistent with the practical codebook selection strategies.
The input to the model consists of $\rho$ sub-array CSI tensors each represented as $\mathcal{H}_{t}\in\mathbb{R}^{N_{x}^{\prime}\times N_{y}^{\prime}\times(N_{r}N_{sc})\times(2N_{pol})}$, where $N_{x}^{\prime}$$N_{y}^{\prime}$ denote the number of horizontal and vertical antenna elements in the sub-array, $N_{r}$ is the number of receive antennas, $N_{sc}$ is the number of subcarriers configured for CSI in the entire bandwidth, and $N_{pol}$ = 2 denotes the dual polarization. The factor of $2$ corresponds to the real and imaginary components of the complex channel co-efficients.
Each CSI tensor is independently processed by a shared encoder module ($\mathrm{Enc}$) to extract the spatial-frequency features. The input is arranged to explicitly preserve the sub-array structure $(N^{\prime}_{x}\times N^{\prime}_{y})$, enabling the convolutional layers to effectively learn channel variations across the sub-panel dimensions. The encoder consists of three consecutive 3D convolutional layers with filter sizes {16,32,64} followed by a Global Average Pooling operation to capture the effective features. It is finally passed through a Dense layer of $D=32$ units, yielding a compact latent representation
$\mathbf{z}_{t}=\mathrm{Enc}(\mathcal{H}_{t}),\;\mathbf{z}_{t}\in\mathbb{R}^{D}$ (4)
To exploit the temporal structure captured by the sequential sub-array sounding, the latent vectors $\{\mathbf{z}_{t}\}_{t=1}^{\rho}$ are passed through Gated Recurrent Unit (GRU).
The model architecture of CsiAdaNet is illustrated in the Fig. 3, employs multiple heads each corresponding to a distinct CSI feedback quantity. All of the heads operate on a shared latent representation obtained after temporal aggregation from the GRU layers and feature refinement through Head Block units. Each prediction is implemented using a lightweight feed-forward network comprising of a Dense layer followed by a softmax activation, thereby formulating the outputs as a classification over the respective discrete levels.
Beam Set Head: It predicts the over-sampled beamset index ($\mathcal{B}$), which is a coarse representation of the spatial direction and rotation in the beam space. The head operates on the shared representation and outputs a probability distribution over the $O_{1}\times O_{2}$ beamset levels.
Beam Indices Head: It predicts the top $L$ beam indices ($\mathcal{I}$) among the $N_{x}\times N_{y}$ orthogonal beams within the selected beamset. It takes in the beamset information in its latent form as an additional information to predict the beam indices accurately.
Wideband Amplitude Head: It predicts the quantized wideband amplitude levels associated with the selected top $L$ beams in the previous stage. The amplitudes capture the power distribution among the selected $2L$ beams and are drawn from a pre-defined finite set $P$.
Wideband Phase: It predicts the quantized wideband phase levels corresponding to the selected beams. The phase capture a way to coherently combine the selected beams to focus energy in the desired direction. The latent beam information from beamset and beam indices prediction are used in predicting the quantized phase levels.
The CSIAdaNet model is trained using a weighted sum of Sparse-Categorical Cross-Entropy (SCCE) losses for individual CSI quantities. For a single sample, the loss is defined as
[…]
where $\hat{y}$ denotes the ground-truth class index and $\mathit{p}_{x}$ is the predicted class probability score. The task specific loss is given by
$\mathcal{L}_{x}=\mathbb{E}[\mathit{l}_{SCCE}(\hat{x},\mathit{p}_{x})],\;x\in\{\mathcal{B},\mathcal{I},\mathcal{P},\mathcal{\varphi}\}$ (6)
The overall training loss is defined as
The predicted codebook can be reconstructed as

### v2

@@ Abstract
Future wireless systems are expected to employ a substantially larger number of transmit ports for channel state information (CSI) estimation compared to current specifications. Although this improves spectral efficiency, it also increases the resource overhead to transmit reference signals across the time-frequency grid, ultimately reducing achievable data throughput. In this work, we propose an deep learning (DL)-based CSI reconstruction framework that serves as an enabler for reliable CSI acquisition in future 6G systems. The proposed solution involves designing a port-cycling mechanism that sequentially sounds different portions of CSI ports across time, thereby lowering the overhead while preserving channel observability. The proposed CSI Adaptive Network (CsiAdaNet) model exploits the resulting sparse measurements and captures both spatial and temporal correlations to accurately reconstruct the full-port CSI. The simulation results show that our method achieves overhead reduction while maintaining high CSI reconstruction accuracy.
@@ I Introduction
Massive multiple-input multiple-output (MIMO) system is a crucial technology for wireless communication systems, offering benefits including spatial multiplexing, capacity gain, etc. These benefits rely on accurate CSI, whose acquisition incurs substantial overhead with increasing channel dimensions. In systems operating in frequency division duplexing (FDD) scenario, downlink channel acquisition is performed using CSI-RS [1], which enables reporting of channel quality, number of independent streams, and the beamforming vector [2]. However, CSI resource overhead grows linearly with the array size, reducing data throughput.
To address the growing overhead, industry and academic efforts have explored Artificial Intelligence/ Machine Learning (AI/ML) techniques for CSI feedback. In particular, 3GPP has identified CSI compression and prediction as initial use-cases for AI/ML based air interface [3]. In academia, CsiNet demonstrated effective non-linear CSI compression [4], while subsequent works investigate CSI extrapolation across time, frequency and spatial domains [5]. However these approaches primarily reduce feedback or temporal overhead and do not address the instantaneous overhead present due to large scale antenna sounding.
@@
Building on this framework, we propose CsiAdaNet, a deep learning model that exploits the temporal and spatial correlations across the temporally aggregated CSI measurements to predict the quantized Type-II precoder defined in 5G NR [1],[7],enabling precoder estimation without requiring full-port instantaneous CSI.
The main contributions of this work are summarized as follows:
  1. 1.
    A port-cycling based CSI sounding technique that reduces the instantaneous spatial overhead in massive MIMO systems.
  2. 2.
    A deep learning model that directly predicts Type-II precoder quantities from temporally aggregated port-cycled measurements.
  3. 3.
    Link-level simulation in a 3GPP compliant setup demonstrate proposed approach achieves near optimal performance.
@@ II System Model
where $\mathbf{y}_{k}\in\mathbb{C}^{N_{r}\times 1},\mathbf{h}_{k}\in\mathbb{C}^{N_{r}\times N_{t}},\mathbf{v}_{k}\in\mathbb{C}^{N_{L}\times N_{t}},\mathbf{x}_{k}\in\mathbb{C}^{N_{L}\times 1},n_{k}$ denotes the received signal, channel matrix, precoding vector, data vector and the additive noise of the $k^{th}$ subcarrier. The BS is equipped with $N_{t}=2\times N_{x}\times N_{y}$ antennas, where $N_{x},N_{y}$ denotes the number of dual polarized antenna elements at the horizontal and vertical dimensions. The wideband covariance matrix is computed as
$\mathbf{R}=\frac{1}{K}\sum_{i=1}^{k}h^{H}_{i}h_{i},\quad\mathbf{R}\in\mathbb{C}^{N_{t}\times N_{t}}$ (2)
@@ II-A Type II codebook in 5G NR
  • Wideband Amplitude: It corresponds to reporting the quantized wideband amplitudes for the selected 2L12L-1 orthogonal beams. The quantized amplitudes are drawn from the set, $P={\{0,\sqrt{\frac{1}{64}},\sqrt{\frac{1}{32}},\sqrt{\frac{1}{16}},\sqrt{\frac{1}{8}},\sqrt{\frac{1}{4}},\sqrt{\frac{1}{2}},1\}}$
@@
1:Input: $\{\mathbf{V}_{b}\}_{b=1}^{B}$, $\mathbf{R}$, $\mathbf{L}$
2:Output: $B^{\star},\mathcal{I}^{\star}$
3:for $b=1$ to $B$ do
4:  $\mathbf{r}\!\leftarrow\!\mathbf{V}_{b}^{H}(\mathbf{R}_{11}+\mathbf{R}_{22})\mathbf{V}_{b}$
5:  $\mathcal{I}_{b}\!\leftarrow\!\operatorname{Top}L(|\operatorname{diag}({r})|)$
6:  $\eta_{b}\!\leftarrow\!\sum_{i\in\mathcal{I}_{b}}|{r}_{i}|$
7:end for
@@
9:return $B^{\star},\mathcal{I}^{\star}$
@@
1:Input: $b^{\star},\mathcal{I}^{\star},\mathbf{R}$Output: $\mathcal{P}^{\star}_{\nu},{\varphi}^{\star}_{\nu}$
@@
3:for $l=1$ to $L$ do
4:  $p^{\nu}_{0,l}\!=\!\mathbf{e}^{H}_{0,\nu}\mathbf{v}_{l},\ p^{\nu}_{1,l}\!=\!\mathbf{e}^{H}_{1,\nu}\mathbf{v}_{l}$
5:end for
6:for $\text{pol}\!\in\!\{0,1\}$ do
7:  $\tilde{\mathbf{a}}^{\nu}_{pol}\!\leftarrow\!|p^{\nu}_{pol}|/\max_{k}|p^{\nu}_{k,pol}|$
8:  ${\theta}_{\nu,pol}\!\leftarrow\!\angle\tilde{\mathbf{a}}^{\nu}_{pol}$
9:  $\mathcal{A}^{\nu}_{pol}\!\leftarrow\!\arg\min_{i}|P_{i}-\tilde{\mathbf{a}}^{\nu}_{pol}|$
10:  ${\Theta}^{\nu}_{pol}\!\leftarrow\!\arg\min_{j}|\Phi_{j}-{\theta}_{\nu,pol}|$
11:end for
@@
13:return $\mathcal{P}^{\star}_{\nu},{\varphi}^{\star}_{\nu}$
@@ II-B Sub-array partitioning
To reduce the number of active ports for sounding, the uniform planar array (UPA) of size ($N_{x}\times N_{y}$) is partitioned into $\rho=\rho_{x}\times\rho_{y}$ non-overlapping sub-panels. Each sub-panel consists of $N^{\prime}_{x}=N_{x}/\rho_{x}$ and $N^{\prime}_{y}=N_{y}/\rho_{y}$, resulting in $N^{\prime}_{t}=N^{\prime}_{x}\times N^{\prime}_{y}\times N_{pol}$ ports per sub-array.
Sub-panel based partitioning, preserves the original antenna spacing, and thereby avoids spatial aliasing in the beam computation, making it well suited for port-cycling based CSI acquisition. An illustrative example of the procedure is shown in Fig. 1.
@@ II-C Port Cycling Scheme
It refers to the sequential activation of non-overlapping sub-arrays across multiple time instances for downlink CSI transmission. At each instant, only one sub-array is sounded and the full antenna array is covered after $\rho$ cycles. This enables instantaneous reduction in spatial overhead while retaining the observability of the full array across time.
The feasibility of this method relies on the temporal coherence of beam related quantities during the port-cycling duration. To quantify validity of this scheme, a variation score ($\mathcal{M}$) is defined as
Figure 2: Plot of variation scores
@@
Fig. 2, illustrates that both beamset and dominant beam index exhibit low variation across port-cycling instances, particularly at moderate-to-high SNR region, validating the feasibility of temporally aggregating port-cycled CSI measurements to predict precoder.
III CsiAdaNet
This section presents a deep learning assisted solution for port cycling based system, where the proposed CSIAdaNet resides at the user equipment (UE) side and operates on $\rho$ periodic CSI sub-panel measurements, to predict the quantized Type-II codebook parameters.
$\{B^{\star},\;\mathcal{I}^{\star},\;\mathcal{P}^{\star},\mathcal{\varphi}^{\star}\}=\mathcal{F}(\{\mathcal{H}_{t+iP}\}_{i=0}^{\rho-1};\theta)$ (4)
where $\mathcal{F}(\cdot)$ denotes the DL model, with parameters $\theta$ and $\mathcal{H}_{t+iP}$ denotes the CSI measurements corresponding to sub-array sounded at time $t+iP$.
The model performs the predictions hierarchically with contextual dependency. The beamset is first predicted, followed by the top $L$ beam indices conditioned on the beamset, and finally the quantized wideband amplitude and phase levels.
The input to the model consists of $\rho$ sub-array CSI tensors, $\mathcal{H}_{t}\in\mathbb{R}^{N_{x}^{\prime}\times N_{y}^{\prime}\times(N_{r}N_{sc})\times(2N_{pol})}$, where the factor of $2$ corresponds to the real and imaginary components of the complex channel co-efficients.
Each CSI tensor is processed by a shared encoder module $\mathrm{Enc}(\cdot)$ to extract the spatial-frequency features while preserving the sub-array structure $(N^{\prime}_{x}\times N^{\prime}_{y})$. It comprises of three 3D convolutional layers with filter sizes {16,32,64} followed by a global average pooling and a dense layer of $D=32$ units, yielding a latent representation $\mathbf{z}_{t}=\mathrm{Enc}(\mathcal{H}_{t})\in\mathbb{R}^{D}$. The sequence $\{\mathbf{z}_{t}\}_{t=1}^{\rho}$ is further processed by a Gated Recurrent Unit (GRU) to capture temporal correlations across port-cycling instances and further feature refinement through Head Block units. The aggregated representation is then fed to multiple task-specific heads, each implemented as a feed-forward classifier operating on the shared latent features.The overall model architecture of CsiAdaNet is illustrated in Fig. 3.
The Beam Set Head predicts the over-sampled beamset index ($\mathcal{B}$), providing a coarse spatial representation . The head operates on the shared representation and outputs a probability distribution over the $O_{1}\times O_{2}$ beamset levels. Conditioned on this, the Beam Indices Head predicts the top $L$ beam indices ($\mathcal{I}$) within the selected beamset. The Wideband Amplitude Head predicts the quantized wideband amplitude levels for the selected $2L$ beams from a pre-defined set $P$, while the Wideband Phase Head predicts the quantized phase levels from $\Phi$ for coherent beam combination.
The CSIAdaNet is trained using a weighted multi-task sparse categorical cross-Entropy (SCCE) loss. For each CSI quantity ($x\in\{\mathcal{B},\mathcal{I},\mathcal{P},\mathcal{\varphi}\}$), the task specific loss is defined as $\mathcal{L}_{x}=\mathbb{E}[\mathit{l}_{SCCE}(\hat{x},\mathit{p}_{x})]$, where $\hat{x}$ denotes the ground-truth class index and $\mathit{p}_{x}$ is the predicted class probability score. The overall training loss is defined as
Finally, the predicted codebook can be reconstructed as